Use logarithmic differentiation to prove D9: d/dx (1/u^n) =( -n/u^n+1)(du/dx)

1
Expert's answer
2022-05-20T04:38:00-0400

"f(x)=\frac{1}{u^n}"

"\ln{f(x)}=\ln{\frac{1}{u^n}}=-n\ln u"

"\frac{d}{dx}\ln{f(x)}=\frac{d}{dx}(-n\ln u)"

"\frac{1}{f(x)}\frac{df(x)}{dx}=-n\frac{1}{u}\frac{du}{dx}"

"\frac{df(x)}{dx}=-nf(x)\frac{1}{u}\frac{du}{dx}=-n\frac{1}{u^{n+1}}\frac{du}{dx}"

"\frac{d}{dx}\frac{1}{u^n}=-n\frac{1}{u^{n+1}}\frac{du}{dx}"